# Stereo-GS: Multi-View Stereo Vision Model for Generalizable 3D Gaussian Splatting Reconstruction

### v1

@@ Abstract.
Generalizable 3D Gaussian Splatting reconstruction showcases advanced Image-to-3D content creation but requires substantial computational resources and large datasets, posing challenges to training models from scratch. Current methods usually entangle the prediction of 3D Gaussian geometry and appearance, which rely heavily on data-driven priors and result in slow regression speeds. To address this, we propose Stereo-GS, a disentangled framework for efficient 3D Gaussian prediction. Our method extracts features from local image pairs using a stereo vision backbone and fuses them via global attention blocks. Dedicated point and Gaussian prediction heads generate multi-view point-maps for geometry and Gaussian features for appearance, combined as GS-maps to represent the 3DGS object. A refinement network enhances these GS-maps for high-quality reconstruction. Unlike existing methods that depend on camera parameters, our approach achieves pose-free 3D reconstruction, improving robustness and practicality. By reducing resource demands while maintaining high-quality outputs, Stereo-GS provides an efficient, scalable solution for real-world 3D content generation. Project page: https://kevinhuangxf.github.io/stereo-gs.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXX12conference: the 33rd ACM International Conference on Multimedia; October 27-31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. We propose Stereo-GS for efficient training and inference of generalizable 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) reconstruction for Image-to-3D generation. Given a single input image, we use a diffusion model to generate multi-view images. Our Stereo-GS then employs a novel multi-view stereo vision model to extract features from these images, predicting 3DGS geometry (point-maps) and appearance in a disentangled manner to construct multi-view GS-maps as a representation of 3DGS objects. Unlike existing methods (Hong et al., 2023; Tang et al., 2024a) such as LGM (Tang et al., 2024a), which produce degraded results with artifacts under limited resources, our approach generates high-quality 3DGS content even with constrained training resources.
[…]
@@ 1. Introduction
Images-to-3D has become a crucial step in 3D content generation, with numerous studies focused on enhancing the quality of generated content (Hong et al., 2023; Szymanowicz et al., 2024; Zou et al., 2024; Tang et al., 2024a). However, the high demand for training resources poses a challenge for creators seeking to train models from scratch. How can we reduce these training resource demands while still maintaining high-quality generation?
@@
The major training resources for Image-to-3D approaches (Hong et al., 2023; Tang et al., 2024a; Xu et al., 2024b) consist of computing power and datasets. To maintain high generation quality, current approaches typically rely on substantial training resources, including large-scale datasets and high-end GPUs, and resulting in significant amounts of training time. When training resources are insufficient, current approaches often experience a severe drop in performance. For example, with fewer training resources, LRM (Hong et al., 2023) can only render images at a resolution of $128\times 128$ size. LGM (Tang et al., 2024a) also suffers from blurry rendering and coarse geometry results due to the time-consuming regression.
@@
The primary reason for this dilemma stems from the current methods’ entanglement of 3DGS geometry and appearance predictions (Kerbl et al., 2023; Tang et al., 2024a; Xu et al., 2024b), which forces models to rely heavily on data-driven priors to achieve generalizability. These priors, learned from large-scale datasets and extensive computing resources, enable 3D reconstruction from sparse or single images. However, as shown in Figure 1, limited training resources lead to significant performance degradation in 3D reconstruction models. This entanglement forces models to depend heavily on data-driven priors for generalization, which become unreliable under resource constraints. Consequently, decoupling geometry-appearance predictions to reduce reliance on data-driven priors is critical to improving training efficiency and robustness.
@@
To reduce the data-driven reliance, a more reasonable way is to consider a disentangled framework that separates the 3D Gaussian prediction into two distinct tasks: the 3D reconstruction of point clouds for geometry and the prediction of Gaussian features for appearance. This disentangled framework allows us to integrate pre-trained deep neural networks as geometry priors (Williams et al., 2019) to accelerate 3D reconstruction training. Subsequently, the appearance parameter optimization benefits significantly from the accurate geometric conditions already established, enabling more efficient convergence. For example, Triplane-Gaussian (Zou et al., 2024) uses Point-E (Nichol et al., 2022) to generate 4096-point clouds from a single image, generating as the geometry prior. These point clouds are then decoded into Gaussian features via a tri-plane decoder to model appearance. However, Point-E (Nichol et al., 2022) can only generate sparse point clouds ($4096$ points) from a single image and result in undermined 3D reconstruction results. To avoid sparse point clouds prediction, we explore stereo vision models to predict dense point clouds for improved reconstruction quality.
@@
Recent stereo vision models (Wang et al., 2024; Leroy et al., 2024) showcase outstanding geometry reconstruction performance from multiple perspective images. Particularly, DUSt3R (Wang et al., 2024) can generate dense 3D point clouds with precise geometry by building pixel-wise 2D-to-3D mappings, and MASt3R (Leroy et al., 2024) can further predict matching features for finding correspondences across multiple images. The 3D point clouds and matching features can serve as the geometry and matching priors to accelerate the 3D reconstruction model training. However, the 3D point clouds from multi-view images can not be directly fused together, since these stereo vision models (Wang et al., 2024; Leroy et al., 2024) are only generalizable in the local pair-wise views. It relies on per-scene optimization to align the point clouds globally, which is time-consuming and impractical for 3D reconstruction in a feed-forward manner.
@@
We propose Stereo-GS to extend these stereo vision models (Wang et al., 2024; Leroy et al., 2024) from local stereo setups into a multi-view stereo framework for generalizable 3D Gaussian splatting (3DGS) reconstruction (Kerbl et al., 2023; Tang et al., 2024a). As shown in Figure 2, our method employs a stereo vision model as a backbone to extract features from local image pairs. These features are then fused using multi-view global attention blocks, enabling the disentangled generation of 3DGS geometry and appearance. Based on the fused multi-view stereo tokens, we design a point prediction head to predict point-maps for geometry and a Gaussian prediction head to predict Gaussian features for appearance. The point-maps and Gaussian features are then combined as the multi-view GS-maps for representing the 3DGS objects. Such a disentangled design for 3DGS geometry and appearance ensures high efficiency in training time.
@@
Current 3D Gaussian reconstruction models (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024) often rely on camera parameters, which can limit their applicability in real-world scenarios. In contrast, our method leverages the unconstrained stereo vision models (Wang et al., 2024; Leroy et al., 2024) to achieve generalizable pose-free 3D Gaussian reconstruction, ensuring high efficiency in inference time and enhancing practicality for real-world applications. In summary, our contributions are listed below:
@@
    We propose Stereo-GS, a novel method that extends stereo vision models (Wang et al., 2024; Leroy et al., 2024) from local pairwise setups to a multi-view stereo framework. By leveraging multi-view global attention blocks with point and Gaussian prediction heads, our method can generate precise 3D point-maps and Gaussian features in a feed-forward manner.
@@
    Our method achieves pose-free 3D Gaussian reconstruction by utilizing unconstrained stereo feature backbones. This ensures robustness and practicality for real-world applications where camera data may be unavailable or unreliable.
@@ 2. Related work
Novel View Synthesis (NVS) aims to generate new, unseen perspectives of an object or scene from a provided set of images, by creating 3D representations (Mildenhall et al., 2019). Neural Radiance Fields (Mildenhall et al., 2021; Luo et al., 2025; Huang et al., 2024a) (NeRF) have achieved photo-realistic representations of 3D scenes, encoded by multilayer perception (MLP) networks (Mildenhall et al., 2021; Müller et al., 2022; Barron et al., 2022). Recently, 3D Gaussian Splatting (Kerbl et al., 2023; Song et al., 2024a; Huang et al., 2024c) (3DGS) has demonstrated remarkable results by rendering each point in the 3D space via an efficient rasterization process. NVS methods (Mildenhall et al., 2021; Kerbl et al., 2023; Song et al., 2024b; Luo et al., 2023) usually require scene-specific optimization, resulting in increased computational demands and longer processing times. Compared with per-scene optimization methods, feed-forward-based methods (Szymanowicz et al., 2024; Wang et al., 2021; Chen et al., 2024b) can achieve generalizable 3D reconstruction on the entire scene in a single pass, eliminating the need for additional optimization.
@@
With advancements in the radiance field, there is a growing interest in generalizable 3D reconstruction from multi-view images (Ni et al., 2024; Yu et al., 2021; Shi et al., 2024), which may contain limited overlaps. Specifically, unconstrained 3D reconstruction methods (Ye et al., 2024; Xu et al., 2024a; Wang et al., 2023; Lin et al., 2024) are attracting significant interest due to their generalization and pose-free capabilities. Recent unconstrained stereo vision frameworks DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024) showcase outstanding 3D reconstruction performance from multi-view images. DUSt3R (Wang et al., 2024) can generate 3D point clouds from pair-wise images by building the 2D-to-3D correspondences. MASt3R (Leroy et al., 2024) can further predict matching features between the pair-wise views. However, both DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024) are only generalizable in local pair-wise views and require post-hoc alignment to merge multi-view point clouds. This motivates us to lift the stereo models from a local pair-wise setting into a multi-view stereo framework.
@@
2D diffusion models (Rombach et al., 2022; Saharia et al., 2022) are initially designed for generating single-view images without the ability for novel view synthesis. The recent progress on multi-view diffusion model (Wang and Shi, 2023; Shi et al., 2023b; Xu et al., 2024c; Huang et al., 2024b) empowers the 2D diffusion model to generate multi-view perspectives for 3D objects, by leveraging the camera poses as the conditional input. However, the Multi-view diffusion model can usually generate only a few views with less than 6 images, and inconsistencies may still occur across the generated views. Recently, video diffusion models (Chen et al., 2024a; Voleti et al., 2025) have demonstrated the capability for object-centric multi-view generation, producing more output frames and achieving even more consistent results. Leveraging the power of multi-view diffusion models or video diffusion models (Chen et al., 2024a; Voleti et al., 2025), we can accomplish Image-to-3D and Text-to-3D tasks. These models generate multi-view images, which serve as inputs for our proposed method to reconstruct 3D objects.
@@ 3. Preliminary of 3DGS
Starting from a sparse set of Structure-from-Motion (SfM) (Snavely et al., 2006) points, the goal of 3DGS (Kerbl et al., 2023) is to optimize a scene representation that enables high-quality novel view synthesis. The scene is modeled as a collection of 3D Gaussians, which represent the radiance emitted in the 3D space around each point. Each 3D Gaussian is parameterized by its mean $\mu\in\mathbb{R}^{3}$ as the position, opacity $\alpha\in\mathbb{R}^{3}$ for transparency, Gaussian covariance matrix $\Sigma$ as she shape, and spherical harmonics $c\in\mathbb{R}^{3\times d}$ as the view-dependent color. By utilizing a scaling matrix $S$ and rotation matrix $R$, we can determine the corresponding $\Sigma=RSS^{T}R^{T}$ and ensure $\Sigma$ is positive semi-definite. In our experiments, we focus on view-independent RGB color $S\in\mathbb{R}^{3}$ in the spherical harmonics for each Gaussian. The 3D Gaussians need to be further projected to 2D Gaussians for rendering by volume splatting (Zwicker et al., 2001) method. During rendering, 3DGS follows a typical neural point-based approach (Kopanas et al., 2022) to compute the color $C$ of a pixel by blending $\mathcal{N}$ depth ordered points:
@@
where $c_{i}$ is the color estimated by the spherical harmonics (SH) coefficients of each Gaussian, and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma^{{}^{\prime}}$ (Yifan et al., 2019) multiplied with a per-point opacity.
@@ 4. Proposed method
We introduce a novel Images-to-3D framework, as illustrated in Figure 2. First, we leverage a pre-trained diffusion model (Chen et al., 2024a) to generate multi-view images from a single input image. Then, we employ our proposed Stereo-GS model to estimate multi-view point-maps for geometry and Gaussian features for appearance. Finally, these outputs are combined as GS-maps and enhanced by a refinement network, ensuring the generation of high-quality 3D Gaussians that accurately represent the 3D objects.
@@ 4.1. Stereo-GS
Current unconstrained stereo vision models, such as DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024), are typically designed for local image pairs, while rely on a time-consuming fine-tuning based global alignment for multi-view stereo vision. To address this limitation, we propose a novel method to lift these stereo vision models from a two-view setting into a multi-view stereo setting, enabling efficient and accurate 3D reconstruction in a feed-forward manner.
@@
Given overlapped local image pairs, stereo vision models like DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024) can predict 3D point maps in local camera coordinates. It first encodes both images with a Siamese ViT (Dosovitskiy et al., 2021) encoder $\operatorname{Enc}$, which yields two token representations $T^{1}$ and $T^{2}$. Then, a pair of twined ViT decoders $\operatorname{Dec^{1}},\operatorname{Dec^{2}}$ sequentially performs cross-attention to the tokens with spatial information:
@@
where $\operatorname{Dec}_{i}^{v}\left(T^{1},T^{2}\right)$ denotes the $i$-th block in local input view $v\in\{1,2\}$ with the input tokens $T^{1},T^{2}$, and $i=1,\ldots,K$ for a decoder with $K$ blocks and initialized with encoder tokens $T_{0}^{1}:=T^{1}$ and $T_{0}^{2}:=T^{2}$. Finally, two Dense Prediction Transformer (DPT) (Ranftl et al., 2021) heads predict the final pixel-wise point-maps $P$. However, these stereo point-maps are in a local camera coordinate system, with only the possibility of post-processing alignment to merge the multiple local stereo point-maps.
Thus, we propose to utilize the stereo vision model as the local backbone to extract stereo feature tokens and reconstruct 3D objects within a unified world coordinate in a feed-forward manner. Unlike current methods that rely on a separated DPT head (Ranftl et al., 2021) for predicting individual point-maps for each view, we utilize every stereo pair tokens $\{T^{1}_{i},T^{2}_{i}\}$ as local features, and then directly concatenating them as multi-view stereo tokens to a multi-view attention head for the global fusion of 3D point-maps in a feed-forward manner:
@@
Current methods (Wang et al., 2024; Leroy et al., 2024) apply separated DPT heads on the learned tokens for each input view image. We propose multi-view self-attention blocks to learn global representation based on the multi-view stereo tokens $\tilde{T}$. We select tokens at the layer of $i=\{3,6,9,12\}$ to resemble features at four different stages. These tokens are then fed into the feature fusion blocks (Ranftl et al., 2021) in the DPT head. The selected tokens $\tilde{T}_{i}$ are subsequently processed by multi-view self-attention blocks:
@@
We incorporate a point prediction head to generate multi-view point-maps $P$, which represent the geometry of 3DGS objects. This head is designed similarly to the DPT head (Ranftl et al., 2021), utilizing four feature fusion blocks that progressively upsample the multi-view tokens $\tilde{T}_{i}$ by a factor of two at each stage. The output layer consists of three convolutional layers: the first upsamples the feature predictions to match the input image resolution, while the last reduces the feature dimensions to three channels for point-map prediction. The final point-map representation is initially at half the input image resolution and is then interpolated to match the input image size.

### v2

@@ Abstract.
Generalizable 3D Gaussian Splatting reconstruction showcases advanced Image-to-3D content creation but requires substantial computational resources and large datasets, posing challenges to training models from scratch. Current methods usually entangle the prediction of 3D Gaussian geometry and appearance, which rely heavily on data-driven priors and result in slow regression speeds. To address this, we propose Stereo-GS, a disentangled framework for efficient 3D Gaussian prediction. Our method extracts features from local image pairs using a stereo vision backbone and fuses them via global attention blocks. Dedicated point and Gaussian prediction heads generate multi-view point-maps for geometry and Gaussian features for appearance, combined as GS-maps to represent the 3DGS object. Unlike existing methods that require camera parameters, our approach reconstructs 3D Gaussians from images without camera dependency. A subsequent refinement network enhances the predicted GS-maps, optionally conditioned on camera poses. By reducing resource demands while maintaining high-quality outputs, Stereo-GS provides an efficient, scalable solution for real-world 3D content generation. Project page: https://kevinhuangxf.github.io/stereo-gs.
Image-to-3D generation; Multi-view stereo; 3D Gaussian Splatting
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755151isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Multimedia content creation
Figure 1. We propose Stereo-GS for efficient training and inference of generalizable 3D Gaussian Splatting (3DGS) (kerbl3Dgaussians) reconstruction for Image-to-3D generation. Given a single input image, we use a diffusion model to generate multi-view images. Our Stereo-GS then employs a novel multi-view stereo vision model to extract features from these images, predicting 3DGS geometry (point-maps) and appearance in a disentangled manner to construct multi-view GS-maps as a representation of 3DGS objects. Unlike existing methods (hong2023lrm; tang2024lgm) such as LGM (tang2024lgm), which produce degraded results with artifacts under limited resources, our approach generates high-quality 3DGS content even with constrained training resources.
@@ 1. Introduction
Images-to-3D has become a crucial step in 3D content generation, with numerous studies focused on enhancing the quality of generated content (hong2023lrm; szymanowicz2024splatter; zou2024triplane; tang2024lgm). However, the high demand for training resources poses a challenge for creators seeking to train models from scratch. How can we reduce these training resource demands while still maintaining high-quality generation?
@@
The major training resources for Image-to-3D approaches (hong2023lrm; tang2024lgm; xu2024grm) consist of computing power and datasets. To maintain high generation quality, current approaches typically rely on substantial training resources, including large-scale datasets and high-end GPUs, and resulting in significant amounts of training time. When training resources are insufficient, current approaches often experience a severe drop in performance. For example, with fewer training resources, LRM (hong2023lrm) can only render images at a resolution of $128\times 128$ size. LGM (tang2024lgm) also suffers from blurry rendering and coarse geometry results due to the time-consuming regression.
@@
The primary reason for this dilemma stems from the current methods’ entanglement of 3DGS geometry and appearance predictions (kerbl3Dgaussians; tang2024lgm; xu2024grm), which forces models to rely heavily on data-driven priors to achieve generalizability. These priors, learned from large-scale datasets and extensive computing resources, enable 3D reconstruction from sparse or single images. However, as shown in Figure 1, limited training resources lead to significant performance degradation in 3D reconstruction models. This entanglement forces models to depend heavily on data-driven priors for generalization, which become unreliable under resource constraints. Consequently, decoupling geometry-appearance predictions to reduce reliance on data-driven priors is critical to improving training efficiency and robustness.
@@
To reduce the data-driven reliance, a more reasonable way is to consider a disentangled framework that separates the 3D Gaussian prediction into two distinct tasks: the 3D reconstruction of point clouds for geometry and the prediction of Gaussian features for appearance. This disentangled framework allows us to integrate pre-trained deep neural networks as geometry priors (williams2019deep) to accelerate 3D reconstruction training. Subsequently, the appearance parameter optimization benefits significantly from the accurate geometric conditions already established, enabling more efficient convergence. For example, Triplane-Gaussian (zou2024triplane) uses Point-E (nichol2022point) to generate 4096-point clouds from a single image, generating as the geometry prior. These point clouds are then decoded into Gaussian features via a tri-plane decoder to model appearance. However, Point-E (nichol2022point) can only generate sparse point clouds ($4096$ points) from a single image and result in undermined 3D reconstruction results. To avoid sparse point clouds prediction, we explore stereo vision models to predict dense point clouds for improved reconstruction quality.
@@
Recent stereo vision models (dust3r_cvpr24; mast3r_eccv24) showcase outstanding geometry reconstruction performance from multiple perspective images. Particularly, DUSt3R (dust3r_cvpr24) can generate dense 3D point clouds with precise geometry by building pixel-wise 2D-to-3D mappings, and MASt3R (mast3r_eccv24) can further predict matching features for finding correspondences across multiple images. The 3D point clouds and matching features can serve as the geometry and matching priors to accelerate the 3D reconstruction model training. However, the 3D point clouds from multi-view images can not be directly fused together, since these stereo vision models (dust3r_cvpr24; mast3r_eccv24) are only generalizable in the local pair-wise views. It relies on per-scene optimization to align the point clouds globally, which is time-consuming and impractical for 3D reconstruction in a feed-forward manner.
@@
We propose Stereo-GS to extend these stereo vision models (dust3r_cvpr24; mast3r_eccv24) from local stereo setups into a multi-view stereo framework for generalizable 3D Gaussian splatting (3DGS) reconstruction (kerbl3Dgaussians; tang2024lgm). As shown in Figure 2, our method employs a stereo vision model as a backbone to extract features from local image pairs. These features are then fused using multi-view global attention blocks, enabling the disentangled generation of 3DGS geometry and appearance. Based on the fused multi-view stereo tokens, we design a point prediction head to predict point-maps for geometry and a Gaussian prediction head to predict Gaussian features for appearance. The point-maps and Gaussian features are then combined as the multi-view GS-maps for representing the 3DGS objects. Such a disentangled design for 3DGS geometry and appearance ensures high efficiency in training time.
@@
Current 3D Gaussian reconstruction models (tang2024lgm; xu2024grm; gslrm2024) often rely on camera parameters, which can limit their applicability in real-world scenarios. In contrast, our method leverages the unconstrained stereo vision models (dust3r_cvpr24; mast3r_eccv24) to achieve generalizable 3D Gaussian reconstruction, ensuring high efficiency in inference time and enhancing practicality for real-world applications. In summary, our contributions are listed below:
@@
    We propose Stereo-GS, a novel method that extends stereo vision models (dust3r_cvpr24; mast3r_eccv24) from local pairwise setups to a multi-view stereo framework. By leveraging multi-view global attention blocks with point and Gaussian prediction heads, our method can generate precise 3D point-maps and Gaussian features in a feed-forward manner.
@@
    Our method reconstructs 3D Gaussians without camera dependency by leveraging unconstrained stereo feature backbones. This design ensures robustness and practicality for real-world applications where camera data may be unavailable or unreliable.
@@ 2. Related work
Novel View Synthesis (NVS) aims to generate new, unseen perspectives of an object or scene from a provided set of images, by creating 3D representations (mildenhall2019local). Neural Radiance Fields (mildenhall2021nerf; luo2025nerf; huang2024geometrysticker) (NeRF) have achieved photo-realistic representations of 3D scenes, encoded by multilayer perception (MLP) networks (mildenhall2021nerf; mueller2022instant; barron2022mip). Recently, 3D Gaussian Splatting (kerbl3Dgaussians; song2024geometry; huang2024gaussianmarker; li2024variational) (3DGS) has demonstrated remarkable results by rendering each point in the 3D space via an efficient rasterization process. NVS methods (mildenhall2021nerf; kerbl3Dgaussians; song2024protecting; luo2023copyrnerf) usually require scene-specific optimization, resulting in increased computational demands and longer processing times. Compared with per-scene optimization methods, feed-forward-based methods (szymanowicz2024splatter; wang2021ibrnet; chen2024mvsplat) can achieve generalizable 3D reconstruction on the entire scene in a single pass, eliminating the need for additional optimization.
@@
With advancements in the radiance field, there is a growing interest in generalizable 3D reconstruction from multi-view images (ni2024colnerf; yu2021pixelnerf; shi2024zerorf), which may contain limited overlaps. Specifically, unconstrained 3D reconstruction methods (ye2024no; xu2024sparp; wang2023pf; lin2024relpose++) are attracting significant interest due to their generalization and pose-free capabilities. Recent unconstrained stereo vision frameworks DUSt3R (dust3r_cvpr24) and MASt3R (mast3r_eccv24) showcase outstanding 3D reconstruction performance from multi-view images. DUSt3R (dust3r_cvpr24) can generate 3D point clouds from pair-wise images by building the 2D-to-3D correspondences. MASt3R (mast3r_eccv24) can further predict matching features between the pair-wise views. However, both DUSt3R (dust3r_cvpr24) and MASt3R (mast3r_eccv24) are only generalizable in local pair-wise views and require post-hoc alignment to merge multi-view point clouds. This motivates us to lift the stereo models from a local pair-wise setting into a multi-view stereo framework.
@@
2D diffusion models (rombach2022high; saharia2022photorealistic) are initially designed for generating single-view images without the ability for novel view synthesis. The recent progress on multi-view diffusion model (wang2023imagedream; shi2023mvdream; xu2023dmv3d; huang2024mv) empowers the 2D diffusion model to generate multi-view perspectives for 3D objects, by leveraging the camera poses as the conditional input. However, the Multi-view diffusion model can usually generate only a few views with less than 6 images, and inconsistencies may still occur across the generated views. Recently, video diffusion models (chen2024v3d; voleti2025sv3d) have demonstrated the capability for object-centric multi-view generation, producing more output frames and achieving even more consistent results. Leveraging the power of multi-view diffusion models or video diffusion models (chen2024v3d; voleti2025sv3d), we can accomplish Image-to-3D and Text-to-3D tasks. These models generate multi-view images, which serve as inputs for our proposed method to reconstruct 3D objects.
@@ 3. Preliminary of 3DGS
Starting from a sparse set of Structure-from-Motion (SfM) (snavely2006photo) points, the goal of 3DGS (kerbl3Dgaussians) is to optimize a scene representation that enables high-quality novel view synthesis. The scene is modeled as a collection of 3D Gaussians, which represent the radiance emitted in the 3D space around each point. Each 3D Gaussian is parameterized by its mean $\mu\in\mathbb{R}^{3}$ as the position, opacity $\alpha\in\mathbb{R}^{3}$ for transparency, Gaussian covariance matrix $\Sigma$ as she shape, and spherical harmonics $c\in\mathbb{R}^{3\times d}$ as the view-dependent color. By utilizing a scaling matrix $S$ and rotation matrix $R$, we can determine the corresponding $\Sigma=RSS^{T}R^{T}$ and ensure $\Sigma$ is positive semi-definite. In our experiments, we focus on view-independent RGB color $S\in\mathbb{R}^{3}$ in the spherical harmonics for each Gaussian. The 3D Gaussians need to be further projected to 2D Gaussians for rendering by volume splatting (zwicker2001ewa) method. During rendering, 3DGS follows a typical neural point-based approach (kopanas2022neural) to compute the color $C$ of a pixel by blending $\mathcal{N}$ depth ordered points:
@@
where $c_{i}$ is the color estimated by the spherical harmonics (SH) coefficients of each Gaussian, and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma^{{}^{\prime}}$ (yifan2019differentiable) multiplied with a per-point opacity.
@@ 4. Proposed method
We introduce a novel Images-to-3D framework, as illustrated in Figure 2. First, we leverage a pre-trained diffusion model (chen2024v3d) to generate multi-view images from a single input image. Then, we employ our proposed Stereo-GS model to estimate multi-view point-maps for geometry and Gaussian features for appearance. Finally, these outputs are combined as GS-maps and enhanced by a refinement network, ensuring the generation of high-quality 3D Gaussians that accurately represent the 3D objects.
@@ 4.1. Stereo-GS
Current unconstrained stereo vision models, such as DUSt3R (dust3r_cvpr24) and MASt3R (mast3r_eccv24), are typically designed for local image pairs, while rely on a time-consuming fine-tuning based global alignment for multi-view stereo vision. To address this limitation, we propose a novel method to lift these stereo vision models from a two-view setting into a multi-view stereo setting, enabling efficient and accurate 3D reconstruction in a feed-forward manner.
@@
Given overlapped local image pairs, stereo vision models like DUSt3R (dust3r_cvpr24) and MASt3R (mast3r_eccv24) can predict 3D point maps in local camera coordinates. It first encodes both images with a Siamese ViT (dosovitskiy2020vit) encoder $\operatorname{Enc}$, which yields two token representations $T^{1}$ and $T^{2}$. Then, a pair of twined ViT decoders $\operatorname{Dec^{1}},\operatorname{Dec^{2}}$ sequentially performs cross-attention to the tokens with spatial information:
@@
where $\operatorname{Dec}_{i}^{v}\left(T^{1},T^{2}\right)$ denotes the $i$-th block in local input view $v\in\{1,2\}$ with the input tokens $T^{1},T^{2}$, and $i=1,\ldots,K$ for a decoder with $K$ blocks and initialized with encoder tokens $T_{0}^{1}:=T^{1}$ and $T_{0}^{2}:=T^{2}$. Finally, two Dense Prediction Transformer (DPT) (Ranftl2021dpt) heads predict the final pixel-wise point-maps $P$. However, these stereo point-maps are in a local camera coordinate system, with only the possibility of post-processing alignment to merge the multiple local stereo point-maps.
Thus, we propose to utilize the stereo vision model as the local backbone to extract stereo feature tokens and reconstruct 3D objects within a unified world coordinate in a feed-forward manner. Unlike current methods that rely on a separated DPT head (Ranftl2021dpt) for predicting individual point-maps for each view, we utilize every stereo pair tokens $\{T^{1}_{i},T^{2}_{i}\}$ as local features, and then directly concatenating them as multi-view stereo tokens to a multi-view attention head for the global fusion of 3D point-maps in a feed-forward manner:
@@
Current methods (dust3r_cvpr24; mast3r_eccv24) apply separated DPT heads on the learned tokens for each input view image. We propose multi-view self-attention blocks to learn global representation based on the multi-view stereo tokens $\tilde{T}$. We select tokens at the layer of $i=\{3,6,9,12\}$ to resemble features at four different stages. These tokens are then fed into the feature fusion blocks (Ranftl2021dpt) in the DPT head. The selected tokens $\tilde{T}_{i}$ are subsequently processed by multi-view self-attention blocks:
@@
We incorporate a point prediction head to generate multi-view point-maps $P$, which represent the geometry of 3DGS objects. This head is designed similarly to the DPT head (Ranftl2021dpt), utilizing four feature fusion blocks that progressively upsample the multi-view tokens $\tilde{T}_{i}$ by a factor of two at each stage. The output layer consists of three convolutional layers: the first upsamples the feature predictions to match the input image resolution, while the last reduces the feature dimensions to three channels for point-map prediction. The final point-map representation is initially at half the input image resolution and is then interpolated to match the input image size.